OpenStudy (anonymous):

If a teacher can finish a study sheet 3 times quicker than the student who takes 40 minutes to finish, what is the equation to determine how many minutes it would take for both teacher and student to finish studysheet if working together?

OpenStudy (shiraz14):

In 1 minute, the student finishes 1/40 of the study sheet; the teacher finishes 3/40 of the study sheet. Working together, they will finish 4/40 (or 0.1) of the study sheet (i.e. they will take 10 minutes to finish the study sheet if working together).

OpenStudy (anonymous):

my choices are 1/x + 1/3x=1 or 1/x +3/x=1 or 1/40 +1/120=1/x

hero (hero):

If one person can do a job in \(m\) hours and another person can do a job in \(n\) hours. Then if they work together, they can do the job in \(\dfrac{m \times n}{m + n}\) hours

OpenStudy (shiraz14):

@Hero: You're correct